Use the black point (plus symbol) to indicate the equilibrium price of a tonne of oranges and the equilibrium quantity of ora2300 Domestic Demand Domestic Supply 2100 No Trade Equilibrium Consumer Surplus PRICE (Dollars per tonne) Producer Surplus 0million. (Hint: Take note of the units on the axes of Based on the previous graph, total surplus in the absence of internatio2300 Domestic Demand Domestic Supply Consumer Surplus Producer Surplus PRICE (Dollars per tonne) World Price 0 400 40 80 120When Zambia allows free trade in oranges, the price of a tonne of oranges in Zambia will be $700. At this price, the quantity
Answer #1

Answer:

Before trade, the equilibrium is determined by domestic demand and supply. The equilibrium price is $1300 and quantity is 200 tons of soybeans

CS is the region above the price line and below the demand. PS is the region below price and above supply curve

2300 Domestic Demand Domestic Supply 2100 PRICE (Dollars per tonne) 300 0 + + 40 80120 160 200 240 280 320 360 QUANTITY (Thou

Total surplus = CS + PS = 0.5*(2300 - 300)*200 = $200,000

With trade, producer surplus is increased and consumer surplus is reduced

2300 Domestic Demand Domestic Supply Consumer Surplus Producer Surplus PRICE (Dollars per tonne) World Price 0 400 40 80 120

At this price of 700, 320 tons are demanded and 80 tons are supplied. Hence it will export (320 - 80) = 240 tons(in thousand)

CS before trade = 0.5*(2300 - 1300)*200 = 100,000

PS before trade = 0.5*(1300 - 300)*200 = 100,000

Now after trade we see that

CS before trade = 0.5*(2300 - 700)*320 = 256,000

PS before trade = 0.5*(700 - 300)*80 = 16,000

CS increases by (1) 256,000 – 100,000 = 156,000 and PS decreases by (100,000 – 16,000) = 84,000. Net effect of trade is that total surplus is gain by (156,000 - 84,000) = 72,000